OpenStudy (anonymous):

A person stands on level ground 60m from the nearest point of a cylindrical tank of radius 20m. Calculate the percentage of the circumference seen.

OpenStudy (dumbcow):

 From diagram, the arc seen is 2x so percentage of circumference is 2x/360 or x/180 \[x = \cos^{-1} (\frac{1}{4})\] \[p = \frac{\cos^{-1} (\frac{1}{4})}{180}\]